Use the following information to answer questions:

Risk factors for obesity. A case control study was conducted to determine the risk factors associated with obesity in young adults. The study included 5,763 cases. Of the cases 62% ate fast food at least 3 times a week. It was a matched case control study and 2 controls were matched per 1 case. Of the controls 54% ate fast food at least 3 times a week.

Using the above data and the measurement you calculated, analyze the result of your calculation and explain what it means.
Answer #1

SOL)

Given data

Number of cases=5763

Number of people in case study = 1/3 * 5763

                                                    = 0.33*5763

                                                    =1921

Number of cases in the control study = 2/3 * 5763

                                                            =0.66*5763

                                                             =3842

Hence, here the 2x2 table

Fast Food study

Yes

NO

Total

Case

1191

730

1921

Control

2075

1767

3842

Total

3266

2497

5763

Here

We know the odd ratio = (ad)/ bc

                                     =(1191 * 1767)/ (2075 * 730)

                                     =2104497/1514750

                                       =1.389

Here We get the odds ratio = 1.389

So, here we can see that there is high 1.389 times more obesity in young adults who take at least 3 times per week than the control.

Here we should now see that we will make a correlation between cases and control study